OpenStudy (anonymous):

Can someone please help me I tried to start it and came up with 7? Please help what is the slope of a line that passes through the point (-5 3) and is parallel to a line that passes through (2,13) and (-4,-11)

OpenStudy (campbell_st):

ok... so the slope of the parallel line is \[m = \frac{13 - (-11)}{2 - (-4)}\] just calculate it out

OpenStudy (campbell_st):

Parallel lines have the same slope... if you want the equation of the line use y = mx + b m is the slope you just found, then substitute x = -5 and y = 3 to find b then you'll have the equation of the line

OpenStudy (anonymous):

How do I find b?

OpenStudy (campbell_st):

ok... what did you get for m..?

OpenStudy (anonymous):

I got 7

OpenStudy (campbell_st):

nope look at this \[m = \frac{13 - (-11)}{2 - (-4)} = \frac{24}{6} =?\]

OpenStudy (anonymous):

4 sorry

OpenStudy (anonymous):

Do I put it in the y-y1=m(X1-x) ?

OpenStudy (campbell_st):

great so the equation of the line is \[y = 4x + b\] substitute x = -5 and y = 3 so \[3 = 4 \times -5 + b\] what is the value of b?

OpenStudy (campbell_st):

you could use point slope form... but I always find slope intercept is easier

OpenStudy (anonymous):

So I multiply 4 & -5?

OpenStudy (campbell_st):

well if you substitute 3 = -20 + b so b = 23 then the equation is y = 4x + 23

OpenStudy (anonymous):

Omg I still don't get it :( thank you for helping me though

OpenStudy (anonymous):

If it's -20 + 3 wouldn't it be 17?